OpenStudy (anonymous):

Use Euler's formula to find the missing number. Faces: Edges:15 Vertices:9

OpenStudy (zzr0ck3r):

So we are talking about Eulers formula for planar graphs. Which states that for a planar graph we have \[v-e+f=2\] So in this situation \(9-15+f=2\). Solve for \(f\).

OpenStudy (rhr12):

F = 2 - V + E

OpenStudy (rhr12):

8 is the answer as i think

OpenStudy (anonymous):

Thank you @rhr12

OpenStudy (rhr12):

You are most welcome